### v1

# Dictionary-Learning-Based Data Pruning for System Identification

Abstract
System identification is normally involved in augmenting time series data by time shifting and nonlinearisation (via polynomial basis), which introduce redundancy both feature-wise and sample-wise. Many research works focus on reducing redundancy feature-wise, while less attention is paid to sample-wise redundancy. This paper proposes a novel data pruning method, called (mini-batch) FastCan, to reduce sample-wise redundancy based on dictionary learning. Time series data is represented by some representative samples, called atoms, via dictionary learning. The useful samples are selected based on their correlation with the atoms. The method is tested on one simulated dataset and two benchmark datasets. The R-squared between the coefficients of models trained on the full and the coefficients of models trained on pruned datasets is adopted to evaluate the performance of data pruning methods. It is found that the proposed method significantly outperforms the random pruning method.
$\dagger$\$$\dagger$\$footnotetext: Equal contribution
1 Introduction
System identification refers to a method of identifying the mathematical description of a dynamic system using measured input-output data [1]. It can be used to forecast future values, assess the effects of input variations, design control schemes, etc. According to the objective of system identification, it can be generally divided into two types [2]. The first type focuses on the approximation scheme that produces the minimum production errors, such as fuzzy logic [3] and neural networks [4]. The second type focuses on the elucidation of the underlying rule that represents the system, such as spectral analysis [5], the Volterra series [6] and Nonlinear Auto-Regressive with eXogenous Inputs (NARX) [7]. NARX-based methods have gained increasing attention in various fields such as engineering, finance, biology, and social sciences due to their flexibility in modelling a variety of systems while maintaining interpretability [8, 2, 9].
To avoid model overfitting and reduce computational complexity, feature selection methods based on orthogonalisation techniques and greedy search are widely applied in constructing NARX models [10, 11, 12]. The orthogonalisation-based method was firstly derived in [10] to efficiently decide which terms should be included in the Nonlinear Autoregressive Moving Average with eXogenous input (NARMAX) models. Then this method was further developed in [11] as orthogonal forward-regression estimators to identify parsimonious models of structure-unknown systems by modifying and augmenting some orthogonal least squares methods. A more comprehensive review of this feature selection idea and its development for system identification can be found in [12].
Similar to feature selection, input selection is also a crucial step in system identification, which can help to improve the identification performance and reduce storage and training costs [13, 14, 12]. While optimal input selection (often referred to as data pruning) is currently an active research topic for computer vision [15, 16, 17] and natural language processing [18, 19], its application to system identification has received less attention, although it remains an intractable task.
This paper proposes a data-pruning method based on dictionary learning to select useful time series samples for constructing the simplest mathematical model to represent a system. Specifically, the reduced polynomial NARX method is applied to the identification of nonlinear dynamic systems as an example. The canonical-correlation-based fast feature selection method introduced in [20] is adopted to find the most important terms that should be included in the NAXR model to achieve the required accuracy. Dictionary learning is then creatively combined with the fast feature selection method based on canonical correlation to find the most useful time series samples to train these terms.
Dictionary learning, initially popular in the signal processing area, is used to learn an overcomplete dictionary composed of signal atoms, where a sparse linear combination of these atoms can generate the original signal [21]. It is currently widely studied for image processing since images usually admit sparse representation [22, 23]. For the proposed method, k-means-based dictionary learning provides an overcomplete dictionary of time terms, which are used as pseudo-labels in the data pruning process.
In this paper, the proposed method, i.e., (mini-batch) FastCan, is compared with the random pruning method. The reduced NARX trained with the full dataset is used as the baseline model. The coefficients learned with the pruned dataset is compared with the baseline model. The more the coefficients are close to the baseline model, the better the pruned dataset is. The procedural of the research in this paper is illustrated in Fig. 1.
The next section introduces the details of the proposed data-pruning method for system identification, followed by numerical case studies to demonstrate the necessity and advantages of the method. Subsequently, case studies on two benchmark datasets for nonlinear system identification are presented, with a discussion on the effect of hyperparamteres in the proposed method. It is worth noting that the proposed method can also be used to select informative samples for general machine-learning tasks.
Figure 1: The schematic diagram of this research. The time series $y(t)$ is generated by the target equation, then time-shifted and nonlinearised to construct a redundant term library. Feature-wise redundancy is represented by a correlation heatmap, while sample-wise redundancy is visualised by a principle component analysis (PCA) plot. Feature selection is applied to the term library to reduce feature-wise redundancy. Dataset A [^1] The colour style used this figure is from bg-mpl-stylesheets [24]., which contains the full set of samples, is used to train the Baseline model. Dataset B, a random subset of Dataset A, is used to train the Random pruned model. Dataset C, consisting of samples selected based on the atoms learned from Dataset A, is used to train the FastCan pruned model. The FastCan pruned dataset generally exhibits less sample-wise redundancy compared to the randomly pruned dataset, as illustrated in the PCA plot, where the data selected by FastCan are more widely spread, while the randomly pruned data tend to be close to each other. To compare the data pruning methods, Dataset A is pruned repeatedly by random selection and FastCan, and the models are then trained on the pruned datasets, respectively. The results show that the coefficients learned from the FastCan pruned data are closer to those of the baseline model, with less variance across repeated tests.
2 Methodology
This section introduces the dictionary learning-based data pruning method for a discrete system with finite orders.
2.1 Reduced polynomial NARX
A NARX model is used to simulate a system, which is formulated as
$\begin{split}y(k)=&F[y(k-1),y(k-2),...,y(k-n_{y}),\\ &u(k-d),y(k-d-1),...,y(k-d-n_{u})]+e(k)\end{split}$ (1)
where $y(k)$, $u(k)$ and $e(k)$ are the system output, input and noise sequences, respectively. $n_{y}$ and $n_{u}$ are the maximum terms for the system output and input. $d$ is the time delay, typically set to $d=1$. $F[\cdot]$ is a nonlinear function.
The power-form polynomial model is adopted to approximate the nonlinear mapping $F[\cdot]$, and the equation (1) is then given as,
$\begin{split}y(k)=\theta_{0}&+\sum_{i_{1}=1}^{n}f_{i_{1}}\big{(}x_{i_{1}}(k)\big{)}\\ &+\sum_{i_{1}=1}^{n}\sum_{i_{2}=i_{1}}^{n}f_{i_{1}i_{2}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k)\big{)}+\dots+\sum_{i_{1}=1}^{n}\cdots\\ &\sum_{i_{\ell}=i_{\ell-1}}^{n}f_{i_{1}i_{2}\cdots i_{\ell}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{\ell}}(k)\big{)}+e(k)\end{split}$ (2)
where $\ell$ is the degree of polynomial nonlinearity, $n=n_{y}+n_{u}$, and
$f_{i_{1}i_{2}\cdots i_{m}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{m}}(k)\big{)}=\theta_{i_{1}i_{2}\cdots i_{m}}\prod_{k=1}^{m}x_{i_{k}}(k),$
$x_{m}(k)=\begin{cases}y(k-m),&1\leq m\leq n_{y},\\ u(k-m+n_{y}),&n_{y}+1\leq m\leq n=n_{y}+n_{u},\end{cases}$
where $1\leq m\leq\ell$, $\theta_{i_{1}i_{2}\cdots i_{m}}$ are model parameters, and $\prod_{k=1}^{m}x_{i_{k}}(k)$ are model terms whose order is not higher than $\ell$.
The total number of model terms in the polynomial NARX model given in equation (2) is $M=(n+\ell)!/[n!\,\ell!]$. It can be seen that the full NARX model can include a large number of terms, increasing the risk of overfitting. In addition, only a subset of these terms is typically important to capture the underlying dynamic relationship [2]. Therefore, the canonical-correlation-based fast feature selection is carried out to find out the significant model terms from $\prod_{k=1}^{m}x_{i_{k}}(k)$ with $y(k)$ as the target. Refer to [20] for details on the specific feature selection steps.
2.2 Mini-batch data pruning
For simplicity, the selected model terms are represented by the matrix $\mathbf{X}$ as follows,
$\mathbf{X}=\begin{pmatrix}x_{1,1}&\dots&x_{1,N}\\ \vdots&\ddots&\vdots\\ x_{p,1}&\dots&x_{p,N}\end{pmatrix}$
where $p$ is the number of the features, that is the selected model terms, and $N$ is the number of time series samples given the selected features.
A dictionary $\mathbf{D}\in\mathbb{R}^{p\times K}$ for the matrix $\mathbf{X}\in\mathbb{R}^{p\times N}$ is obtained by mini-batch k-means clustering [25] and used as the target $\mathbf{Y}$ for the subsequent sample selection step, where $K$ is the number of atoms in the dictionary, and typically $K>p$ for an overcomplete dictionary. Each data sample $\mathbf{x}_{j}\in\mathbb{R}^{p}$ (column of $\mathbf{X}$) can be approximated as a sparse linear combination of the dictionary atoms, given by,
$\mathbf{x}_{j}\approx\mathbf{D}\mathbf{a}_{j}$
where $\mathbf{a}_{j}$ is a sparse coefficient vector for the $j$-th sample, which means most of the entries in $\mathbf{a}_{j}$ are zero.
The fast selection method based on canonical correlation is performed again, with $\mathbf{D}$ as the target, to find $q$ significant samples from the sample matrix $\mathbf{X}$. For selection methods which evaluate the linear association between candidates, no additional information is gained when the number of selected samples $q$ exceeds the rank of the data matrix $p$. Therefore, to avoid invalid sample selection due to $q>p$, samples are selected in separate batches by the mini-batch sample selection algorithm, which is given in Algorithm 1. Within each batch, the redundancy and interaction of samples are considered, while they are ignored between batches.
Input: $X\in\mathbb{R}^{p\times N}$ (sample matrix), $Y\in\mathbb{R}^{p\times K}$ (target matrix), $q$ (number of samples to select)
Output: $d$ (selected indices)
Step 1: Compute the batch size $b=$ $\lceil q/K\rceil$ and then generate the batch matrix $B\in\mathbb{R}^{K\times s}$, where $B[i,j]\leq b$ and $\sum B=q$;
Step 2: Initialize the candidate sample matrix $X_{c}\leftarrow X$;
for $i\leftarrow 1$ to $K$ do
       Let $y_{i}\leftarrow Y[:,i]$;
       for $j\leftarrow 1$ to $s$ do
             for $k\leftarrow 1$ to $B[i,j]$ do
                   Compute the MCC between each column of $X_{c}$ and $y_{i}$;
                   Select the index $l^{*}$ corresponding to the largest MCC;
                   Append $l^{*}$ to $d$;
                   Remove $X_{c}[:,l^{*}]$ from the candidate matrix;
return $d$;
Algorithm 1 Mini-batch Sample Selection
3 Numerical Case Studis
3.1 Input Data Analysis
Data pruning is crucial in system identification for the following reasons:
  1. 1.
    Firstly, system identification typically involves time-series data sampled from continuous physical processes. When the sampling interval is small, consecutive data points are highly similar, leading to a strong temporal correlation and redundancy. The larger dataset size due to redundant data increases the computational cost of model training.
  2. 2.
    Secondly, the time-shift operation, commonly employed in system identification introduces delayed versions of the signals as new features, which makes the redundancy of the dataset worse. Redundant features do not provide additional information but increase the complexity of the model, making it more prone to overfitting and harder to interpret.
[…]
$t$ (s) $y(t)$ $y(t-\Delta t)$ $y(t-2\Delta t)$ $y(t-20\Delta t)$
0 0 0 0 0
0.01 0.063 0 0 0
0.02 0.127 0.063 0 0
0.03 0.189 0.127 0.063 0
0.04 0.251 0.189 0.127 0
0.99 -$0$0.063 -0.127 -0.955
Table 1: Time-shifted data representation with $\Delta t=0.01\,\text{s}$.
The feature-wise redundancy is illustrated in the correlation heatmap of Fig. 3, where the $i^{th}$ feature corresponds to $y(t-i\Delta t)$. The high Pearson’s correlation near the main diagonal indicates redundancy between neighbouring features, such as $y(t-i\Delta t)$ and $y(t-(i+1)\Delta t)$. This kind of redundancy can be mitigated via feature selection.
Fig. 4 shows the redundancy within the time series samples using principal component analysis (PCA). After projecting the 21-dimensional data into two-dimensional space, it is found that the data forms a continuous trajectory over time. The proximity of adjacent samples in the lower-dimensional space given by PCA indicates that there is redundancy between these time series samples. This redundancy can be addressed by data pruning.
Figure 3: Correlation heatmap illustrating feature-wise redundancy
Figure 4: Correlation heatmap illustrating sample-wise redundancy
3.2 Data with Dual Stable Equilibria
The simulated time series data is generated by a non-autonomous nonlinear system, given by,
$\ddot{y}+\dot{y}-y+y^{2}+y^{3}=u$
where $u(t)=0.1\cos(0.2\pi t)$. By setting $u(t)=0$ and initialising $y$ and $\dot{y}$ with different values, a phase portrait of the nonlinear system is obtained, as shown in Fig. 5.
Figure 5: Phase portrait of the nonlinear system with dual stable equilibria
It can be seen that this system exhibits two stable equilibria. To comprehensively capture its dynamics, the data pruning algorithm should identify two distinct types of samples attracted to the different stable spirals and select samples from both categories.
Unlike feature selection, random selection serves as a robust benchmark for sample selection and often outperforms more complex algorithms [26]. Therefore, the results of the proposed data-pruning method will be compared with those of the random selection method to demonstrate its advantages.
The structure of the dual-stable-equilibria (DSE) data is visualized using its first and second principal components. The selected samples and atoms are projected onto these principal directions, with their distribution shown in Fig. 6. It can be found that the samples selected via mini-batch sample selection are distributed across the entire dataset, whereas the randomly selected samples are concentrated primarily on the left and right regions, with less coverage in the middle.
Figure 6: Visualization of the DSE data structure
To quantitatively evaluate the selection performance of the mini-batch and random methods, a reduced polynomial NARX model with 10 terms is derived using the full dataset as the baseline. Note that in all case studies, the number of terms for system output and input is ten respectively, so the total number of terms is $20$, and the degree of polynomial nonlinearity is three. The model coefficients trained on the full dataset are used as the baseline results. The performance of the baseline NARX on the test dataset is provided in Appendix 1. The NARX terms are then fixed, and the model coefficients are trained using the pruned datasets obtained via the mini-batch and random selection methods. The number of atoms for dictionary learning is set to 100, and 600 samples are then selected. The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared. The sample selection and training processes are repeated ten times, with the results presented in Fig. 7.
The mini-batch method gives results with a higher median and lower variance, whereas the random selection method performs significantly worse. The same phenomenon is observed when the number of selected samples varies between 300 and 600, as illustrated in Fig. 8. Furthermore, for the mini-batch method, the variance of the model coefficients decreases as the number of selected samples increases. However, this relationship does not exist when the samples are randomly selected.
Figure 7: Comparison of sample selection performance between mini-batch and random selection methods on DSE data using R-squared
Figure 8: The effect of sample size on the performance of mini-batch and random selection methods for DSE data
4 Case Studies on the Benchmark Datasets
Two benchmark datasets, collected from the Electro-Mechanical Positioning System (EMPS) [27] and the Wiener-Hammerstein System (WHS) [28] are adopted to evaluate the performance of the mini-batch method for data pruning in real system identification.
4.1 Data from the Electro-Mechanical Positioning System
The structure of the data collected from the EMPS, a standard drive configuration for prismatic joints of robots or machine tools, is visualised by its first two principal components, as shown in Fig. 9. The primary nonlinearity of the corresponding data is introduced by friction effects.
The atoms determined for the dictionary learning and the selected samples are also shown in Fig 9. The samples selected by the mini-batch method are spread over all parts of the dataset, while the randomly selected samples are concentrated in the middle two parts but pay less attention to the left and the right parts.
Figure 9: Visualization of the EMPS data structure
Figure 10: Comparison of sample selection performance between mini-batch and random selection methods on EMPS data using R-squared
The baseline NARX model with 10 terms is derived again, and the model coefficients trained on the full dataset are used as the baseline to quantitatively evaluate the selection performance of the mini-batch and random methods. See Appendix 1 for the performance of the baseline NARX on the test dataset. The number of atoms for dictionary learning is set to 100, and 10000 samples are then selected. The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared, with the NARX terms fixed. The sample selection and training processes are repeated ten times, with the results illustrated in Fig. 10. In addition, Fig. 11 shows the results corresponding to two methods when the number of selected samples varies between 2000 and 20000.
Figure 11: The effect of sample size on the performance of mini-batch and random selection methods for EMPS data
Similar to the observations from the DSE data, the mini-batch method generally produces results with a higher median and lower variance. However, there are exceptions when the number of selected samples is less than 6000, where the randomly selected samples yield higher mean values. This scenario will be analysed and further discussed in Section 4.3. Additionally, as expected, the variance of the model coefficients decreases as the number of selected samples increases.
4.2 Data from the Wiener-Hammerstein System
The structure of the data collected from the Wiener-Hammerstein System (WHS), a well-known block-oriented structure consisting of a static nonlinearity between two linear time-invariant blocks, is visualised in Fig. 12 using its first two principal components. The atoms determined for the dictionary learning and the selected samples are also shown in Fig 14. Different from the observations from the DSE data and the EMPS data, the samples obtained by the mini-batch method and random selection method have similar distributions, with the selected samples distributed over the entire dataset.
The same strategy is applied to quantitatively evaluate the sample selection performance of the mini-batch and random methods for the identification task of WHS. The baseline NARX model with 10 terms is derived, and its performance on the test dataset is provided in Appendix 1. The number of atoms for dictionary learning is set to 100, and 10000 samples are selected, with the R-squared results shown in Fig. 13. Then, the number of selected samples varies between 2000 and 20000 to demonstrate the variation of R-squared with the selected sample size, as presented in Fig. 14.
Figure 12: Visualization of the WHS data structure
Figure 13: Comparison of sample selection performance between mini-batch and random selection methods on WHS data using R-squared
Figure 14: The effect of sample size on the performance of mini-batch and random selection methods for WHS data
Fig. 13 demonstrates that the results obtained by the mini-batch and random methods are similar. This phenomenon is also observed in Fig. 14 when the number of the selected samples exceeds 8000. This result is reasonable and expected, as similar sample distributions provide similar information for system identification. In addition, similar to the EMPS case, when the number of selected samples is small (less than 8000), the randomly selected samples yield higher mean values. This phenomenon will be analysed and further discussed in the following section.
4.3 The Effect of Hyperparameters in the Mini-batch Method
In the mini-batch method, sample selection results are influenced by two hyperparameters: the number of atoms in the dictionary and the batch size. To explore how these two hyperparameters affect the performance of the selected samples, system identification results against the varying atom size and batch size for the three previously used datasets are given in Fig. 15 to Fig. 16. For the DSE dataset, $600$ samples are selected, while for EMPS and WHS datasets, $10000$ samples are selected.
It can be noticed that the batch size in Algorithm 1 is determined by both the number of samples and the number of atoms, with the aim of minimising redundancy among the selected samples when all atoms are used in the sample selection process. Therefore, for the results in Fig. 15, the batch size varies for each atom size. Notably, a larger atom size does not leads to better performance, because some atoms may be learned from noise rather than meaningful data. In Fig. 16, when changing the batch size for each dataset, the number of atoms is fixed at the best atom size obtained from Fig. 15. In this scenario, using a larger batch size to reduce redundancy can improve the performance of the selected samples.
(a) DSE dataset
(b) EMPS dataset
(c) WHS dataset
Figure 15: The effect of atom size on the performance of mini-batch method
(a) DSE dataset
(b) EMPS dataset
(c) WHS dataset
Figure 16: The effect of batch size on the performance of mini-batch method with the best atom size
To investigate whether a larger batch size always leads to better selection when the number of atoms is fixed, the relationship between R-squared and batch size for EMPS data with different atom sizes is presented in Fig. 17. The results indicate that the relationship between R-squared and batch size is not a straightforward linear relationship.
In summary, the optimal atom size is different for each dataset, and the optimal batch size may not be the ratio of the sample number to the atom size. A more sophisticated method is required to determine the appropriate hyperparameters for the proposed method. Furthermore, the possible reason for the previous-mentioned phenomenon shown in Fig. 11 and Fig. 14 is that when the number of selected samples is small, the performance of the selected samples may be more sensitive to the effectiveness of the hyperparameters.
(a) atom size 10
(b) atom size 100
(c) atom size 1000
Figure 17: The effect of batch size on the performance of mini-batch method with difference atom sizes for EMPS data
5 Conclusion
This paper introduces a data-pruning method based on dictionary learning to choose appropriate time-series samples for identification tasks of a discrete system with finite orders. The selection process is independent of the specific system identification technique used.
The case studies on the synthetic and benchmark datasets show that the proposed method effectively reduces the sample size without significantly degrading identification performance. Compared with random selection, the data-pruning method gives better results with a higher mean and lower variance of R-squared values. Additionally, it is found that the careful selection of hyperparamters is crucial, particularly when the number of selected samples is small.
Appendix 1: Prediction Performance of the Baseline NARX
(a) Test 1
(b) Test 2
Figure 18: The performance of the baseline NARX model for DSE test data with different initial conditions
Figure 19: The performance of the baseline NARX model for EMPS test data
Figure 20: The performance of the baseline NARX model for WHS test data

### v2

@@
 Abstract
-System identification is normally involved in augmenting time series data by time shifting and nonlinearisation (via polynomial basis), which introduce redundancy both feature-wise and sample-wise.
+System identification is normally involved in augmenting time series data by time shifting and nonlinearisation (e.g., polynomial basis), both of which introduce redundancy in features and samples.
 Many research works focus on reducing redundancy feature-wise, while less attention is paid to sample-wise redundancy.
-This paper proposes a novel data pruning method, called (mini-batch) FastCan, to reduce sample-wise redundancy based on dictionary learning.
+This paper proposes a novel data pruning method, called mini-batch FastCan, to reduce sample-wise redundancy based on dictionary learning.
 Time series data is represented by some representative samples, called atoms, via dictionary learning.
@@
 The method is tested on one simulated dataset and two benchmark datasets.
-The R-squared between the coefficients of models trained on the full and the coefficients of models trained on pruned datasets is adopted to evaluate the performance of data pruning methods.
+The R-squared between the coefficients of models trained on the full datasets and the coefficients of models trained on pruned datasets is adopted to evaluate the performance of data pruning methods.
 It is found that the proposed method significantly outperforms the random pruning method.
 $\dagger$\$$\dagger$\$footnotetext: Equal contribution
+Keywords Data pruning, system identification, dictionary learning, NARX
 1 Introduction
@@
 To avoid model overfitting and reduce computational complexity, feature selection methods based on orthogonalisation techniques and greedy search are widely applied in constructing NARX models [10, 11, 12].
-The orthogonalisation-based method was firstly derived in [10] to efficiently decide which terms should be included in the Nonlinear Autoregressive Moving Average with eXogenous input (NARMAX) models.
-Then this method was further developed in [11] as orthogonal forward-regression estimators to identify parsimonious models of structure-unknown systems by modifying and augmenting some orthogonal least squares methods.
+The orthogonalisation-based method was firstly derived in [10] to efficiently decide which terms should be included in the Nonlinear Autoregressive Moving Average with eXogenous input models.
+This method was then further developed in [11] as orthogonal forward-regression estimators to identify parsimonious models of structure-unknown systems by modifying and augmenting some orthogonal least squares methods.
 A more comprehensive review of this feature selection idea and its development for system identification can be found in [12].
 Similar to feature selection, input selection is also a crucial step in system identification, which can help to improve the identification performance and reduce storage and training costs [13, 14, 12].
-While optimal input selection (often referred to as data pruning) is currently an active research topic for computer vision [15, 16, 17] and natural language processing [18, 19], its application to system identification has received less attention, although it remains an intractable task.
-This paper proposes a data-pruning method based on dictionary learning to select useful time series samples for constructing the simplest mathematical model to represent a system.
+While optimal selection of input samples (often referred to as data pruning) is currently an active research topic for computer vision [15, 16, 17] and natural language processing [18, 19], its application to system identification has received less attention, although it remains an intractable task.
+Data pruning plays a vital role in system identification for the following reasons.
+First, time-series data collected from continuous physical processes often exhibit strong temporal correlation, especially when sampled at high frequencies, resulting in redundant observations that hinder efficient model training.
+Second, the common use of time-shift operations to generate delayed signal versions further exacerbates data redundancy.
+This paper proposes a data-pruning method based on dictionary learning to select useful time series samples for constructing a concise mathematical model to represent a system accurately.
 Specifically, the reduced polynomial NARX method is applied to the identification of nonlinear dynamic systems as an example.
-The canonical-correlation-based fast feature selection method introduced in [20] is adopted to find the most important terms that should be included in the NAXR model to achieve the required accuracy.
-Dictionary learning is then creatively combined with the fast feature selection method based on canonical correlation to find the most useful time series samples to train these terms.
+It is worth noting that the proposed method can also be used to select informative samples for other types of system identification models.
+The canonical-correlation-based fast feature selection method introduced in [20, 21] is adopted to find the most important terms that should be included in the NAXR model to achieve the required accuracy.
+Dictionary learning is then creatively combined with the Fast selection method based on Canonical correlation (FastCan) to find the most useful time series samples to train these terms.
 Dictionary learning, initially popular in the signal processing area, is used to learn an overcomplete dictionary composed of signal atoms, where a sparse linear combination of these atoms can generate the original signal [22].
 It is currently widely studied for image processing since images usually admit sparse representation [23, 24].
-For the proposed method, k-means-based dictionary learning provides an overcomplete dictionary of time terms, which are used as pseudo-labels in the data pruning process.
-In this paper, the proposed method, i.e., (mini-batch) FastCan, is compared with the random pruning method.
+For the proposed method, k-means-based dictionary learning, introduced in Section 2.2, provides a dictionary of time terms, which is used as pseudo-labels in the data pruning process.
+In this paper, the proposed method, named mini-batch FastCan, is compared with the random pruning method.
+Random pruning commonly serves as a reliable benchmark in sample selection tasks and often yields better results than more advanced methods when 30% or less of the data is retained [25].
 The reduced NARX trained with the full dataset is used as the baseline model.
-The coefficients learned with the pruned dataset is compared with the baseline model.
+The coefficients learned with the pruned dataset is compared with the coefficients of the baseline model.
 The more the coefficients are close to the baseline model, the better the pruned dataset is.
-The procedural of the research in this paper is illustrated in Fig.
+The schematic diagram of the research in this paper is illustrated in Fig.
 1.
 The next section introduces the details of the proposed data-pruning method for system identification, followed by numerical case studies to demonstrate the necessity and advantages of the method.
-Subsequently, case studies on two benchmark datasets for nonlinear system identification are presented, with a discussion on the effect of hyperparamteres in the proposed method.
-It is worth noting that the proposed method can also be used to select informative samples for general machine-learning tasks.
+Subsequently, case studies on two public benchmark datasets [26, 27] for nonlinear system identification are presented, with a discussion on the effect of hyperparamteres in the proposed method.
 Figure 1: The schematic diagram of this research.
 The time series $y(t)$ is generated by the target equation, then time-shifted and nonlinearised to construct a redundant term library.
-Feature-wise redundancy is represented by a correlation heatmap, while sample-wise redundancy is visualised by a principle component analysis (PCA) plot.
+Feature-wise redundancy is represented by a correlation heatmap, while sample-wise redundancy is visualised by a principal component analysis (PCA) plot.
 Feature selection is applied to the term library to reduce feature-wise redundancy.
@@
 Dataset B, a random subset of Dataset A, is used to train the Random pruned model.
-Dataset C, consisting of samples selected based on the atoms learned from Dataset A, is used to train the FastCan pruned model.
-The FastCan pruned dataset generally exhibits less sample-wise redundancy compared to the randomly pruned dataset, as illustrated in the PCA plot, where the data selected by FastCan are more widely spread, while the randomly pruned data tend to be close to each other.
-To compare the data pruning methods, Dataset A is pruned repeatedly by random selection and FastCan, and the models are then trained on the pruned datasets, respectively.
-The results show that the coefficients learned from the FastCan pruned data are closer to those of the baseline model, with less variance across repeated tests.
+Dataset C, consisting of samples selected based on the atoms learned from Dataset A, is used to train the Mini-batch FastCan pruned model.
+The selected samples obtained by both methods are visualised via a PCA plot.
+To compare the data pruning methods quantitatively, Dataset A is pruned repeatedly by random selection and mini-batch FastCan, and the models are then trained on the pruned datasets, respectively.
+Data pruning performance is then evaluated by the R-squared score between the coefficients learned from the pruned samples and the coefficients of the baseline model.
 2 Methodology
-This section introduces the dictionary learning-based data pruning method for a discrete system with finite orders.
+This section introduces the dictionary learning-based data pruning method, mini-batch FastCan, for a discrete system with finite orders.
 2.1 Reduced Polynomial NARX
-A NARX model is used to simulate a system, which is formulated as
-$\begin{split}y(k)=&F[y(k-1),y(k-2),...,y(k-n_{y}),\\ &u(k-d),y(k-d-1),...,y(k-d-n_{u})]+e(k)\end{split}$ (1)
+The reduced polynomial NARX model is a nonlinear dynamic model that represents the system output as a sparse subset of polynomial terms consisting of past outputs and inputs selected from the full NARX model structure.
+It reduces complexity by including only important or relevant terms, rather than all possible combinations.
+Mathematically, a NARX model is used to simulate a system, which is formulated as
+$\begin{split}y(k)=&F[y(k-1),y(k-2),...,y(k-n_{y}),\\ &u(k-1),u(k-2),...,u(k-n_{u})]+e(k)\end{split}$ (1)
 where $y(k)$, $u(k)$ and $e(k)$ are the system output, input and noise sequences, respectively.
-$n_{y}$ and $n_{u}$ are the maximum terms for the system output and input.
-$d$ is the time delay, typically set to $d=1$.
+$n_{y}$ and $n_{u}$ are the maximum lags for the system output and input.
 $F[\cdot]$ is a nonlinear function.
 The power-form polynomial model is adopted to approximate the nonlinear mapping $F[\cdot]$, and the equation (1) is then given as,
-$\begin{split}y(k)=\theta_{0}&+\sum_{i_{1}=1}^{n}f_{i_{1}}\big{(}x_{i_{1}}(k)\big{)}\\ &+\sum_{i_{1}=1}^{n}\sum_{i_{2}=i_{1}}^{n}f_{i_{1}i_{2}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k)\big{)}+\dots+\sum_{i_{1}=1}^{n}\cdots\\ &\sum_{i_{\ell}=i_{\ell-1}}^{n}f_{i_{1}i_{2}\cdots i_{\ell}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{\ell}}(k)\big{)}+e(k)\end{split}$ (2)
+$\begin{split}y(k)=\theta_{0}&+\sum_{i_{1}=1}^{n}f_{i_{1}}\big{(}x_{i_{1}}(k)\big{)}\\ &+\sum_{i_{1}=1}^{n}\sum_{i_{2}=i_{1}}^{n}f_{i_{1}i_{2}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k)\big{)}+\dots\\ &+\sum_{i_{1}=1}^{n}\cdots\sum_{i_{\ell}=i_{\ell-1}}^{n}f_{i_{1}i_{2}\cdots i_{\ell}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{\ell}}(k)\big{)}+e(k)\end{split}$ (2)
 where $\ell$ is the degree of polynomial nonlinearity, $n=n_{y}+n_{u}$, and
-$f_{i_{1}i_{2}\cdots i_{m}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{m}}(k)\big{)}=\theta_{i_{1}i_{2}\cdots i_{m}}\prod_{k=1}^{m}x_{i_{k}}(k),$
-$x_{m}(k)=\begin{cases}y(k-m),&1\leq m\leq n_{y},\\ u(k-m+n_{y}),&n_{y}+1\leq m\leq n=n_{y}+n_{u},\end{cases}$
-where $1\leq m\leq\ell$, $\theta_{i_{1}i_{2}\cdots i_{m}}$ are model parameters, and $\prod_{k=1}^{m}x_{i_{k}}(k)$ are model terms whose order is not higher than $\ell$.
+$f_{i_{1}i_{2}\cdots i_{d}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{d}}(k)\big{)}=\theta_{i_{1}i_{2}\cdots i_{d}}\prod_{j=1}^{d}x_{i_{j}}(k),\quad$ $1\leq d\leq\ell$
+$x_{i_{j}}(k)=\begin{cases}y(k-i),&1\leq i\leq n_{y}\\ u(k-i+n_{y}),&n_{y}+1\leq i\leq n=n_{y}+n_{u}\end{cases}$
+where $\theta_{i_{1}i_{2}\cdots i_{d}}$ are model parameters, and $\prod_{j=1}^{d}x_{i_{j}}(k)$ are model terms whose order is not higher than $\ell$.
 The total number of model terms in the polynomial NARX model given in equation (2) is $M=(n+\ell)!/[n!\,\ell!]$.
 It can be seen that the full NARX model can include a large number of terms, increasing the risk of overfitting.
-In addition, only a subset of these terms is typically important to capture the underlying dynamic relationship [2].
-Therefore, the canonical-correlation-based fast feature selection is carried out to find out the significant model terms from $\prod_{k=1}^{m}x_{i_{k}}(k)$ with $y(k)$ as the target.
-Refer to [20] for details on the specific feature selection steps.
-2.2 Mini-batch data pruning
+Because only a subset of these terms is typically important to capture the underlying dynamic relationship [2], the canonical-correlation-based fast feature selection is carried out here to find out the $m$ significant model terms from $\prod_{j=1}^{d}x_{i_{j}}(k)$ with $y(k)$ as the target.
+Therefore, the reduced polynomial NARX model can be given as,
+$y(k)=\theta_{0}+\sum_{j=1}^{m}\sum_{i_{1}=1}^{n}\cdots\sum_{i_{d}=i_{d-1}}^{n}f_{i_{1}i_{2}\cdots i_{d}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{d}}(k)\big{)}+e(k),\quad 1\leq m\leq M$
+Refer to [20, 21] for details on the specific feature selection steps.
+2.2 Data Pruning with Mini-batch FastCan
 For simplicity, the selected model terms are represented by the matrix $\mathbf{X}$ as follows,
-$\mathbf{X}=\begin{pmatrix}x_{1,1}&\dots&x_{1,N}\\ \vdots&\ddots&\vdots\\ x_{p,1}&\dots&x_{p,N}\end{pmatrix}$
-where $p$ is the number of the features, that is the selected model terms, and $N$ is the number of time series samples given the selected features.
-A dictionary $\mathbf{D}\in\mathbb{R}^{p\times K}$ for the matrix $\mathbf{X}\in\mathbb{R}^{p\times N}$ is obtained by mini-batch k-means clustering [25] and used as the target $\mathbf{Y}$ for the subsequent sample selection step, where $K$ is the number of atoms in the dictionary, and typically $K>p$ for an overcomplete dictionary.
-Each data sample $\mathbf{x}_{j}\in\mathbb{R}^{p}$ (column of $\mathbf{X}$) can be approximated as a sparse linear combination of the dictionary atoms, given by,
+$\mathbf{X}=\begin{pmatrix}x_{1,1}&\dots&x_{1,N}\\ \vdots&\ddots&\vdots\\ x_{m,1}&\dots&x_{m,N}\end{pmatrix}$
+where $m$ is the number of the features, i.e., the selected model terms, and $N$ is the number of time series samples given these selected features.
+A dictionary $\mathbf{D}\in\mathbb{R}^{m\times q}$ for the matrix $\mathbf{X}\in\mathbb{R}^{m\times N}$ is obtained by mini-batch k-means clustering [29] and used as the target matrix for the subsequent data pruning step, where $q$ is the number of atoms in the dictionary, and typically $q>m$ for an overcomplete dictionary.
+Each data sample $\mathbf{x}_{j}\in\mathbb{R}^{m}$ (column of $\mathbf{X}$) can be approximated as a sparse linear combination of the dictionary atoms, given by,
 $\mathbf{x}_{j}\approx\mathbf{D}\mathbf{a}_{j}$
-where $\mathbf{a}_{j}$ is a sparse coefficient vector for the $j$-th sample, which means most of the entries in $\mathbf{a}_{j}$ are zero.
-The fast selection method based on canonical correlation is performed again, with $\mathbf{D}$ as the target, to find $q$ significant samples from the sample matrix $\mathbf{X}$.
-For selection methods which evaluate the linear association between candidates, no additional information is gained when the number of selected samples $q$ exceeds the rank of the data matrix $p$.
-Therefore, to avoid invalid sample selection due to $q>p$, samples are selected in separate batches by the mini-batch sample selection algorithm, which is given in Algorithm 1.
-Within each batch, the redundancy and interaction of samples are considered, while they are ignored between batches.
-Input: $X\in\mathbb{R}^{p\times N}$ (sample matrix), $Y\in\mathbb{R}^{p\times K}$ (target matrix), $q$ (number of samples to select)
-Output: $d$ (selected indices)
-Step 1: Compute the batch size $b=$ $\lceil q/K\rceil$ and then generate the batch matrix $B\in\mathbb{R}^{K\times s}$, where $B[i,j]\leq b$ and $\sum B=q$;
-Step 2: Initialize the candidate sample matrix $X_{c}\leftarrow X$;
-for $i\leftarrow 1$ to $K$ do
-       Let $y_{i}\leftarrow Y[:,i]$;
-       for $j\leftarrow 1$ to $s$ do
-             for $k\leftarrow 1$ to $B[i,j]$ do
-                   Compute the MCC between each column of $X_{c}$ and $y_{i}$;
-                   Select the index $l^{*}$ corresponding to the largest MCC;
-                   Append $l^{*}$ to $d$;
-                   Remove $X_{c}[:,l^{*}]$ from the candidate matrix;
-return $d$;
-Algorithm 1 Mini-batch Sample Selection
+where $\mathbf{a}_{j}\in\mathbb{R}^{q}$ is a sparse coefficient vector for the $j$-th sample, which means most of the entries in $\mathbf{a}_{j}$ are zero.
+Input: $\mathbf{X}\in\mathbb{R}^{m\times N}$ ;
+  $\triangleright$ Sample matrix
+$q\in\mathbb{N}$ ;
+  $\triangleright$ Number of atoms in a dictionary $\mathbf{D}$
+$p\in\mathbb{N}$ ;
+  $\triangleright$ Batch size; Optional
+$n\in\mathbb{N}$ ;
+  $\triangleright$ Number of samples to select
+Output: $\mathbf{s}\in\mathbb{R}^{1\times n}$ ;
+  $\triangleright$ Selected indices
+Step 1:
+Apply the k-means-based dictionary learning [29] to $\mathbf{X}^{\top}$ with $q$ clusters;
+The resulting $q$ cluster centres form the columns of a dictionary $\mathbf{D}\in\mathbb{R}^{m\times q}$ ;
+  $\triangleright$ Target matrix
+Step 2:
+if $p$ is not specified or $p>\lceil n/q\rceil$ then
+       Set $p\leftarrow\lceil n/q\rceil$;
+      if $p>m$ then
+             Set $p\leftarrow m$;
+Step 3:
+Generate the batch matrix $\mathbf{B}\in\mathbb{R}^{q\times t}$, where $t=$ $\lceil n/(q\times p)\rceil$, $\mathbf{B}[i,j]$ $\leq p$ and $\sum_{i=1}^{q}\sum_{j=1}^{t}\mathbf{B}[i,j]=n$;
+Step 4:
+Initialize the candidate sample matrix $\mathbf{X}_{c}\leftarrow\mathbf{X}$ and the target matrix $\mathbf{D}_{c}\leftarrow\mathbf{D}$;
+for $i\leftarrow 1$ to $q$ do
+       Let $\mathbf{d}_{i}\leftarrow\mathbf{D}_{c}[:,i]$;
+       for $j\leftarrow 1$ to $t$ do
+             Select $\mathbf{B}[i,j]$ samples from $\mathbf{X}_{c}$ by using the canonical-correlation-based fast selection method [21], with $\mathbf{d}_{i}$ serving as the target vector;
+             Append the indices of the selected samples to $\mathbf{s}$;
+             Remove the selected samples from the candidate matrix $\mathbf{X}_{c}$;
+return $\mathbf{s}$;
+Algorithm 1 Mini-batch FastCan
+The canonical-correlation-based fast selection method [21] is performed again, with $\mathbf{D}$ as the target, to find $n$ significant samples from the sample matrix $\mathbf{X}$.
+For selection methods which evaluate the linear association between candidates, no additional information is gained once the number of selected samples $n$ exceeds the rank of the data matrix $m$.
+Therefore, to avoid invalid sample selection due to $n>m$, samples are selected in separate batches by the mini-batch FastCan method, whose pseudocode is given in Algorithm 1, and the corresponding codes are available in the GitHub repository  [^2] https://github.com/MatthewSZhang/data-pruning-sysid [^3] https://github.com/scikit-learn-contrib/fastcan.
+Within each batch, the redundancy and interaction between samples are considered, while these are ignored between batches.
+In Algorithm 1, if the batch size $p$ is not specified, it is set to $\lceil q/k\rceil$.
+If a given $p>\lceil n/q\rceil$, it is also reset to $\lceil q/k\rceil$, since exceeding this threshold will result in some atoms being excluded from the selection process.
+Additionally, this algorithm allows the number of atoms $q$ to be smaller than the number of features $m$, enabling a larger batch size to better capture sample redundancy.
 3 Numerical Case Studis
-3.1 Input Data Analysis
-Data pruning is crucial in system identification for the following reasons:
-  1. 1.
-    Firstly, system identification typically involves time-series data sampled from continuous physical processes.
-When the sampling interval is small, consecutive data points are highly similar, leading to a strong temporal correlation and redundancy.
-The larger dataset size due to redundant data increases the computational cost of model training.
-  2. 2.
-    Secondly, the time-shift operation, commonly employed in system identification introduces delayed versions of the signals as new features, which makes the redundancy of the dataset worse.
-Redundant features do not provide additional information but increase the complexity of the model, making it more prone to overfitting and harder to interpret.
+3.1 Visualisation of Sample Redundancy in System Identification
 […]
 […]
-$t$ (s) $y(t)$ $y(t-\Delta t)$ $y(t-2\Delta t)$ $y(t-20\Delta t)$
-0 0 0 0 0
-0.01 0.063 0 0 0
-0.02 0.127 0.063 0 0
-0.03 0.189 0.127 0.063 0
-0.04 0.251 0.189 0.127 0
-0.99 -$0$0.063 -0.127 -0.955
 Table 1: Time-shifted data representation with $\Delta t=0.01\,\text{s}$.
+$t$ (s) $y(t-\Delta t)$ $y(t-2\Delta t)$ $y(t-3\Delta t)$ $y(t-20\Delta t)$
+0 NaN NaN NaN NaN
+0.01 0.000 NaN NaN NaN
+0.02 0.063 0.000 NaN NaN
+0.03 0.127 0.063 0.000 NaN
+0.04 0.189 0.127 0.063 NaN
+0.99 -0.063 -0.127 -0.189 -0.955
+(a) feature-wise
+(b) sample-wise
+Figure 3: Correlation heatmap illustrating redundancy in data
 The feature-wise redundancy is illustrated in the correlation heatmap of Fig.
-3, where the $i^{th}$ feature corresponds to $y(t-i\Delta t)$.
+3a, where the $i^{th}$ feature corresponds to $y(t-i\Delta t)$ with $1\leq i\leq 20$.
 The high Pearson’s correlation near the main diagonal indicates redundancy between neighbouring features, such as $y(t-i\Delta t)$ and $y(t-(i+1)\Delta t)$.
@@
 Fig.
-4 shows the redundancy within the time series samples using principal component analysis (PCA).
-After projecting the 21-dimensional data into two-dimensional space, it is found that the data forms a continuous trajectory over time.
+3b shows the redundancy within the time series samples using principal component analysis (PCA).
+After projecting the 20-dimensional data into two-dimensional space, it is found that the data forms a continuous trajectory over time.
 The proximity of adjacent samples in the lower-dimensional space given by PCA indicates that there is redundancy between these time series samples.
 This redundancy can be addressed by data pruning.
-Figure 3: Correlation heatmap illustrating feature-wise redundancy
-Figure 4: Correlation heatmap illustrating sample-wise redundancy
 3.2 Data with Dual Stable Equilibria
@@
 where $u(t)=0.1\cos(0.2\pi t)$. By setting $u(t)=0$ and initialising $y$ and $\dot{y}$ with different values, a phase portrait of the nonlinear system is obtained, as shown in Fig. 4.
-Figure 5: Phase portrait of the nonlinear system with dual stable equilibria
-It can be seen that this system exhibits two stable equilibria.
-To comprehensively capture its dynamics, the data pruning algorithm should identify two distinct types of samples attracted to the different stable spirals and select samples from both categories.
-Unlike feature selection, random selection serves as a robust benchmark for sample selection and often outperforms more complex algorithms [26].
+Figure 4: A phase portrait of the nonlinear system with dual stable equilibria
+It can be seen that this system exhibits two stable equilibria, with five measurements for each of the left and right equilibria.
+Therefore, this dataset is referred to as symmetrical dual-stable-equilibria (SDSE) dataset.
+To comprehensively capture its dynamics, the data pruning algorithm should identify two distinct kinds of samples attracted to the different stable spirals and select samples from both kinds.
+Unlike feature selection, random selection serves as a strong baseline for sample selection and often outperforms more sophisticated algorithms when retaining 30% or less of the data [25].
 Therefore, the results of the proposed data-pruning method will be compared with those of the random selection method to demonstrate its advantages.
-The structure of the dual-stable-equilibria (DSE) data is visualized using its first and second principal components.
-The selected samples and atoms are projected onto these principal directions, with their distribution shown in Fig.
-6.
-It can be found that the samples selected via mini-batch sample selection are distributed across the entire dataset, whereas the randomly selected samples are concentrated primarily on the left and right regions, with less coverage in the middle.
-Figure 6: Visualization of the DSE data structure
-To quantitatively evaluate the selection performance of the mini-batch and random methods, a reduced polynomial NARX model with 10 terms is derived using the full dataset as the baseline.
-Note that in all case studies, the number of terms for system output and input is ten respectively, so the total number of terms is $20$, and the degree of polynomial nonlinearity is three.
-The model coefficients trained on the full dataset are used as the baseline results.
-The performance of the baseline NARX on the test dataset is provided in Appendix 1.
-The NARX terms are then fixed, and the model coefficients are trained using the pruned datasets obtained via the mini-batch and random selection methods.
-The number of atoms for dictionary learning is set to 100, and 600 samples are then selected.
-The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared.
+To quantitatively evaluate the selection performance of the mini-batch FastCan and random methods, a reduced polynomial NARX model with ten terms and an intercept is derived using the full dataset as the baseline.
+The model coefficients trained on the full training dataset are used as the baseline results.
+A more detailed description of the baseline NARX model and its prediction performance on the test dataset are provided in Appendix 1.
+The NARX terms are then fixed, and the model coefficients are trained using the pruned training datasets obtained via the mini-batch FastCan and random selection methods.
+100 samples are then selected as an example, and the corresponding number of atoms for dictionary learning is set to 15.
+The reasons for this number of atoms will be discussed in Section 5.
+The coefficients trained on the two pruned datasets are compared with those trained on the full dataset using the R-squared.
 The sample selection and training processes are repeated ten times, with the results presented in Fig.
-7.
-The mini-batch method gives results with a higher median and lower variance, whereas the random selection method performs significantly worse.
-The same phenomenon is observed when the number of selected samples varies between 300 and 600, as illustrated in Fig.
-8.
-Furthermore, for the mini-batch method, the variance of the model coefficients decreases as the number of selected samples increases.
-However, this relationship does not exist when the samples are randomly selected.
-Figure 7: Comparison of sample selection performance between mini-batch and random selection methods on DSE data using R-squared
-Figure 8: The effect of sample size on the performance of mini-batch and random selection methods for DSE data
+5a.
+(a) SDSE dataset
+(b) ADSE dataset
+Figure 5: Comparison of sample selection performance between mini-batch FastCan and random selection methods using R-squared on two datasets
+As shown in Fig.
+5a, the mini-batch FastCan method gives results with a higher median and lower variance, indicating superior and more consistent performance compared to random selection.
+As illustrated in Fig.
+6a, the same pattern is observed when the number of selected samples varies between 50 and 150 and the number of atoms is fixed to 15, except for 60 samples—this case will be discussed in Section 5.
+Furthermore, for the mini-batch FastCan method, the variance of the model coefficients decreases as the number of selected samples increases.
+However, this relationship does not exist when the samples are randomly selected in Fig.
+6a.
+(a) SDSE dataset
+(b) ADSE dataset
+Figure 6: The effect of sample size on the performance of mini-batch FastCan and random selection methods for two datasets. SD stands for standard deviation.
+To examine whether the sample distributions corresponding to the results shown in Fig.
+5a are different, the structure of the SDSE data is visualised using its first and second principal components, as shown in Fig.
+7a.
+The selected samples corresponding to the results shown in Fig.
+5a are projected onto these principal directions, with their distribution shown in Fig.
+7a.
+It can be observed that the samples selected via mini-batch FastCan tend to cluster around the learned atoms, and the randomly selected samples are evenly distributed across the entire dataset according to the spatial distribution of candidate samples.
+Compared to random selection, the proposed method makes more samples distributed at the left and right ends.
+Nevertheless, the overall difference between the two sample distributions is relatively small.
+(a) SDSE dataset
+(b) ADSE dataset
+Figure 7: Visualisation of the data structure of two datasets
+To further determine whether the sample distribution selected by the proposed method is different from that of the random method, 2 measurements in left equilibrium and 98 measurements in right equilibrium are simulated to repeat the previous analysis.
+This dataset is referred to as asymmetrical dual-stable-equilibria (ADSE) dataset.
+100 samples are selected, and the number of atoms is set to 20 (see Section 5 for details on this setting).
+The results are presented in Figs.
+5b-7b.
+As shown in Figs.
+5b and 6b, when the number of samples across different conditions is imbalanced, the mini-batch FastCan method has obvious advantages over the random sample selection method.
+Furthermore, as shown in Fig.
+7b, almost all samples selected by the random method are concentrated in the right area, that is, these samples belong to the right equilibrium.
+In contrast, the proposed method selects samples from both equilibria, with those from the left equilibrium evenly distributed in the left area.
+These phenomena indicate that, unlike random selection, the mini-batch FastCan method is less influenced by the imbalance in the candidate sample distribution and is more effective at capturing diverse system behaviours.
+This phenomenon, in turn, helps explain why the identification performance corresponding to the proposed method is superior to that corresponding to the random selection method.
 4 Case Studies on the Benchmark Datasets
-Two benchmark datasets, collected from the Electro-Mechanical Positioning System (EMPS) [27] and the Wiener-Hammerstein System (WHS) [28] are adopted to evaluate the performance of the mini-batch method for data pruning in real system identification.
+Two benchmark datasets, collected from the Electro-Mechanical Positioning System (EMPS) [26] and the Wiener-Hammerstein System (WHS) [27] are adopted to evaluate the performance of the mini-batch FastCan method for data pruning in real-world system identification tasks.
 4.1 Data from the Electro-Mechanical Positioning System
-The structure of the data collected from the EMPS, a standard drive configuration for prismatic joints of robots or machine tools, is visualised by its first two principal components, as shown in Fig.
-9.
+The EMPS is a standard drive configuration for prismatic joints of robots or machine tools.
 The primary nonlinearity of the corresponding data is introduced by friction effects.
-The atoms determined for the dictionary learning and the selected samples are also shown in Fig 9.
-The samples selected by the mini-batch method are spread over all parts of the dataset, while the randomly selected samples are concentrated in the middle two parts but pay less attention to the left and the right parts.
-Figure 9: Visualization of the EMPS data structure
-Figure 10: Comparison of sample selection performance between mini-batch and random selection methods on EMPS data using R-squared
-The baseline NARX model with 10 terms is derived again, and the model coefficients trained on the full dataset are used as the baseline to quantitatively evaluate the selection performance of the mini-batch and random methods.
+The baseline NARX model with ten terms and an intercept is derived, and the model coefficients trained on the full training dataset are used as the baseline to quantitatively evaluate the selection performance of the mini-batch FastCan and random methods.
 See Appendix 1 for the performance of the baseline NARX on the test dataset.
-The number of atoms for dictionary learning is set to 100, and 10000 samples are then selected.
+The number of selected samples is 100, and the number of atoms for dictionary learning is set to 25 (see Section 5 for details on this setting).
 The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared, with the NARX terms fixed.
 The sample selection and training processes are repeated ten times, with the results illustrated in Fig.
-10.
+8a.
 In addition, Fig.
-11 shows the results corresponding to two methods when the number of selected samples varies between 2000 and 20000.
-Figure 11: The effect of sample size on the performance of mini-batch and random selection methods for EMPS data
-Similar to the observations from the DSE data, the mini-batch method generally produces results with a higher median and lower variance.
-However, there are exceptions when the number of selected samples is less than 6000, where the randomly selected samples yield higher mean values.
-This scenario will be analysed and further discussed in Section 4.3.
-Additionally, as expected, the variance of the model coefficients decreases as the number of selected samples increases.
+9a presents the results corresponding to two methods for different numbers of selected samples, ranging from 20 to 120.
+(a) EMPS dataset
+(b) WHS dataset
+Figure 8: Comparison of sample selection performance between mini-batch FastCan and random selection methods using R-squared on two benchmark datasets
+(a) EMPS dataset
+(b) WHS dataset
+Figure 9: The effect of sample size on the performance of mini-batch FastCan and random selection methods for two benchmark datasets. SD stands for standard deviation.
+Consistent with the observations from the SDSE and ADSE datasets, the mini-batch FastCan method produces results with a higher median and lower variance, indicating more stable and reliable performance.
+Additionally, as expected, the variance of the model coefficients also decreases as the number of selected samples increases for the mini-batch FastCan method.
+The structure of the data collected from the EMPS is visualised by its first two principal components, as shown in Fig.
+10a.
+The atoms determined for the dictionary learning and the selected samples corresponding to the results given in Fig.
+8a are projected onto this reduced space.
+The atoms and selected samples corresponding to the mini-batch FastCan method are distributed in the central, left and right regions, while the randomly selected samples are almost only spread over the central part.
+(a) EMPS dataset
+(b) WHS dataset
+Figure 10: Visualisation of the data structure of two benchmark datasets
 4.2 Data from the Wiener-Hammerstein System
-The structure of the data collected from the Wiener-Hammerstein System (WHS), a well-known block-oriented structure consisting of a static nonlinearity between two linear time-invariant blocks, is visualised in Fig.
-12 using its first two principal components.
-The atoms determined for the dictionary learning and the selected samples are also shown in Fig 14.
-Different from the observations from the DSE data and the EMPS data, the samples obtained by the mini-batch method and random selection method have similar distributions, with the selected samples distributed over the entire dataset.
-The same strategy is applied to quantitatively evaluate the sample selection performance of the mini-batch and random methods for the identification task of WHS.
-The baseline NARX model with 10 terms is derived, and its performance on the test dataset is provided in Appendix 1.
-The number of atoms for dictionary learning is set to 100, and 10000 samples are selected, with the R-squared results shown in Fig.
+The WHS is a well-known block-oriented structure consisting of a static nonlinearity between two linear time-invariant blocks.
+The same strategy is applied to quantitatively evaluate the sample selection performance of the mini-batch FastCan and random methods for the identification task of WHS.
+The baseline NARX model with ten terms and an intercept is derived, and its performance on the test dataset is provided in Appendix 1.
+The number of atoms for dictionary learning is set to 5 (see Section 5 for details on this setting, and 100 samples are selected, with the R-squared results shown in Fig.
+8b.
+Then, the number of selected samples varies between 20 and 120 to demonstrate the variation of R-squared with the selected sample size, as presented in Fig.
+9b.
+Fig.
+8b demonstrates that the results obtained by the mini-batch FastCan are better than those obtained by the random method.
+This phenomenon is also observed in Fig.
+9b except at 40 samples.
+This exception will be analysed and further discussed in the following section.
+The structure of the data collected from the WHS is visualised in Fig.
+10b using its first two principal components.
+The atoms determined for the dictionary learning and the selected samples are shown in Fig.
+10b, with the number of atoms set to 5.
+Different from the observations from the ADSE data and the EMPS data, the samples obtained by the mini-batch FastCan method and the random selection method have similar distributions, with the selected samples mainly distributed over the central part of the dataset.
+This phenomenon is unexpected because the samples associated with the results in Fig.
+8b exhibit similar distributions but provide varying levels of information for system identification.
+This observation suggests that sample selection strategies based solely on distributional characteristics of the data space may be insufficient for identifying informative samples.
+(a) SDSE dataset
+(b) ADSE dataset
+(c) EMPS dataset
+(d) WHS dataset
+Figure 11: The effect of atom size on the performance of mini-batch FastCan method
+5 The Effect of Hyperparameters in the Mini-batch FastCan Method
+In the mini-batch FastCan method, sample selection results are influenced by two hyperparameters: the number of atoms in the dictionary and the batch size.
+To examine how these two hyperparameters affect the performance of the selected samples, system identification results against the varying atom size and batch size for the four previously used datasets are given in Fig.
+11 to Fig.
+12.
+For all datasets, $100$ samples are selected.
+For the results in Fig.
+11, the batch size is not given as an input parameter but is instead set to the ceiling of the ratio between the number of selected samples and the number of atoms ($\lceil n/q\rceil$), with the aim of minimising redundancy among the selected samples when all atoms are used in the sample selection process.
+Therefore, the batch size varies for each atom size in Fig.
+11.
+As shown in Fig.
+11, the performance deteriorates when the atom size exceeds the size associated with the highest R-squared value.
+This may be attributed to the fact that, for a fixed number of selected samples, larger atom sizes lead to smaller batch sizes, thus increasing sample-wise redundancy.
+In practical engineering applications, the optimal atom size can be determined by defining a candidate range and selecting the value that yields the best performance via an exhaustive search.
+(a) SDSE dataset
+(b) ADSE dataset
+(c) EMPS dataset
+(d) WHS dataset
+Figure 12: The effect of batch size on the performance of mini-batch FastCan method under the optimal atom size
+(a) SDSE dataset
+(b) ADSE dataset
+(c) EMPS dataset
+(d) WHS dataset
+Figure 13: The effect of atom size on the performance of mini-batch FastCan method under different sample numbers
+(a) SDSE dataset
+(b) WHS dataset
+Figure 14: Comparison of sample selection performance between the mini-batch FastCan method with tuned atom size and random selection using R-squared on two datasets
+In Fig.
+12, when changing the batch size for each dataset, the number of atoms is fixed at the optimal atom size obtained from Fig.
+11, where the optimal atom sizes $q$ for the SDSE, ADSE, EMPS, and WHS datasets are 15, 20, 25, and 5, respectively.
+According to Step 2 in Algorithm 1, the maximum batch sizes for the DSE, EMPS, and WHS datasets are 7, 5, 4 and 10, respectively.
+It can be seen that when the atom size is fixed at the optimal value, using a larger batch size to reduce redundancy can generally improve the performance of the selected samples.
+Therefore, the atom sizes and the batch sizes for the previous four case studies are the optimal atom sizes obtained from Fig.
+11 and the corresponding maximum batch sizes.
+However, it is worth noting that in Fig.
+12, the relationship between R-squared and batch size is not a straightforward linear relationship, especially for the EMPS and WHS datasets.
+A more sophisticated method can be developed in the future to determine the appropriate hyperparameters for the proposed method.
+To examine the underlying causes of the previously observed exceptions—where random selection outperforms or matches the performance of the mini-batch FastCan method, as shown in Figs.
+6a and 9b—Fig.
+13 presents the impact of atom size on model performance across varying selected-sample sizes for the mini-batch FastCan approach.
+This analysis aims to assess whether suboptimal hyperparameter settings contribute to these discrepancies.
+To investigate the potential causes for the previously noted exceptions where random selection outperforms or matches mini-batch FastCan selection in Figs.
+6a and 9b, the effect of atom size on model performance corresponding to the mini-batch FastCan method across different selected-sample sizes is demonstrated in Fig.
 13.
-Then, the number of selected samples varies between 2000 and 20000 to demonstrate the variation of R-squared with the selected sample size, as presented in Fig.
+For each sample size, the maximum R-squared value is printed out.
+It can be observed that when the mini-batch FastCan method performs close to or worse than the random method, the adopted atom size is not the optimal size for the given number of selected samples.
+For the ADSE dataset, when 60 samples are selected, the optimal atom size is 5 rather than 15.
+For the WHS dataset, when 40 samples are selected, the optimal atom size is 10 rather than 5.
+After tuning the atom size for the mini-batch FastCan method, the comparison of sample selection performance of the two methods is then illustrated in Fig.
 14.
-Figure 12: Visualization of the WHS data structure
-Figure 13: Comparison of sample selection performance between mini-batch and random selection methods on WHS data using R-squared
-Figure 14: The effect of sample size on the performance of mini-batch and random selection methods for WHS data
-Fig.
-13 demonstrates that the results obtained by the mini-batch and random methods are similar.
-This phenomenon is also observed in Fig.
-14 when the number of the selected samples exceeds 8000.
-This result is reasonable and expected, as similar sample distributions provide similar information for system identification.
-In addition, similar to the EMPS case, when the number of selected samples is small (less than 8000), the randomly selected samples yield higher mean values.
-This phenomenon will be analysed and further discussed in the following section.
-4.3 The Effect of Hyperparameters in the Mini-batch Method
-In the mini-batch method, sample selection results are influenced by two hyperparameters: the number of atoms in the dictionary and the batch size.
-To explore how these two hyperparameters affect the performance of the selected samples, system identification results against the varying atom size and batch size for the three previously used datasets are given in Fig.
-15 to Fig.
-16.
-For the DSE dataset, $600$ samples are selected, while for EMPS and WHS datasets, $10000$ samples are selected.
-It can be noticed that the batch size in Algorithm 1 is determined by both the number of samples and the number of atoms, with the aim of minimising redundancy among the selected samples when all atoms are used in the sample selection process.
-Therefore, for the results in Fig.
-15, the batch size varies for each atom size.
-Notably, a larger atom size does not leads to better performance, because some atoms may be learned from noise rather than meaningful data.
-In Fig.
-16, when changing the batch size for each dataset, the number of atoms is fixed at the best atom size obtained from Fig.
-15.
-In this scenario, using a larger batch size to reduce redundancy can improve the performance of the selected samples.
-(a) DSE dataset
-(b) EMPS dataset
-(c) WHS dataset
-Figure 15: The effect of atom size on the performance of mini-batch method
-(a) DSE dataset
-(b) EMPS dataset
-(c) WHS dataset
-Figure 16: The effect of batch size on the performance of mini-batch method with the best atom size
-To investigate whether a larger batch size always leads to better selection when the number of atoms is fixed, the relationship between R-squared and batch size for EMPS data with different atom sizes is presented in Fig.
-17.
-The results indicate that the relationship between R-squared and batch size is not a straightforward linear relationship.
-In summary, the optimal atom size is different for each dataset, and the optimal batch size may not be the ratio of the sample number to the atom size.
-A more sophisticated method is required to determine the appropriate hyperparameters for the proposed method.
-Furthermore, the possible reason for the previous-mentioned phenomenon shown in Fig.
-11 and Fig.
-14 is that when the number of selected samples is small, the performance of the selected samples may be more sensitive to the effectiveness of the hyperparameters.
-(a) atom size 10
-(b) atom size 100
-(c) atom size 1000
-Figure 17: The effect of batch size on the performance of mini-batch method with difference atom sizes for EMPS data
-5 Conclusion
-This paper introduces a data-pruning method based on dictionary learning to choose appropriate time-series samples for identification tasks of a discrete system with finite orders.
-The selection process is independent of the specific system identification technique used.
-The case studies on the synthetic and benchmark datasets show that the proposed method effectively reduces the sample size without significantly degrading identification performance.
-Compared with random selection, the data-pruning method gives better results with a higher mean and lower variance of R-squared values.
-Additionally, it is found that the careful selection of hyperparamters is crucial, particularly when the number of selected samples is small.
+The results indicate that the mini-batch FastCan method can provide better sample selection results when the appropriate atom size is applied.
+In addition, as shown in Figs.
+13a, 13c and 13d, when the number of selected samples exceeds a certain threshold, the optimal atom size no longer changes with the number of selected samples.
+At the same time, Figs.
+6a, 9a and 9b indicate that the performance of the selected samples also stabilises when the number of selected samples exceeds this threshold for each dataset.
+This consistent phenomenon suggests that when more samples cannot contribute additional information, the number of optimal atoms learned from the selected samples tends to stabilise.
+This finding supports, to a certain extent, the rationality and robustness of using dictionary learning to generate pseudo labels.
+Moreover, careful tuning of hyperparameters is particularly important when the number of selected samples is small.
+6 Conclusions
+This paper introduces a dictionary learning-based data pruning method, called mini-batch FastCan, for selecting informative time-series samples in the identification of discrete systems with finite orders.
+Specifically, k-means-based dictionary learning is well-suited for generating pseudo-labels in the data pruning process, mitigating the impact of imbalanced candidate sample distributions on sample selection.
+Additionally, selecting samples in separate batches allows for the selection of more samples than the number of necessary model terms, while effectively accounting for redundancy within each batch.
+This method helps ensure that the selected samples are more diverse and informative.
+The selection process is independent of the specific system identification technique used, which enhances its versatility.
+Although hyperparameters, including atom size and batch size, can be empirically determined by engineering practices, the development of theoretically grounded optimisation strategies is recommended to improve generalizability and methodological rigour in future work.
+The case studies on the synthetic and benchmark datasets show that the proposed method effectively reduces the sample size without significantly degrading the identification performance of the benchmark system.
+Compared to random selection, the data-pruning method generally yields better results, with a higher mean and lower variance of R-squared values.
+Additionally, it is found that the careful selection of hyperparameters is crucial, particularly when the proportion of selected samples is small.
+Acknowledgements
+The authors would like to acknowledge the support for this work from Shanghai Qi Zhi Institute Innovation Program (SQZ202310) and National Natural Science Foundation of China (52378187).
+For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any arising Author Accepted Manuscript version.
 Appendix 1: Prediction Performance of the Baseline NARX
+For each of the four datasets, a reduced polynomial NARX model with ten terms and an intercept is derived using the full dataset as a baseline.
+To select relevant nonlinear terms, the maximum input and output lags are set to 4 for the SDSE system, ADSE system, and EMPS, and to 7 for the WHS.
+For all systems, the polynomial degree is fixed at three, and the number of selected model terms is limited to ten.
+These parameter settings are designed to provide reasonable predictive performance, but can be further optimised based on specific application requirements.
 (a) Test 1
 (b) Test 2
-Figure 18: The performance of the baseline NARX model for DSE test data with different initial conditions
-Figure 19: The performance of the baseline NARX model for EMPS test data
-Figure 20: The performance of the baseline NARX model for WHS test data
+Figure 15: The performance of the baseline NARX model for SDSE test data with different initial conditions
+(a) Test 1
+(b) Test 2
+Figure 16: The performance of the baseline NARX model for ADSE test data with different initial conditions
+(a) EMPS dataset
+(b) WHS dataset
+Figure 17: The performance of the baseline NARX model for two benchmark datasets